OpenStudy (anonymous):

find the area bounded by the two curve x=y^2-1 and x=|y| sqrt(1-y^2)

OpenStudy (anonymous):

I got 2

OpenStudy (anonymous):

You're the guy who helped me! Can you please help me with convergence later? lol

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

\[2\int\limits_{0}^{1}y \sqrt{1-y^2}dy-2\int\limits_{0}^{1}(y^2-1)dy\]

OpenStudy (anonymous):

use substitution for the first integral

OpenStudy (anonymous):

u=1-y^2

OpenStudy (anonymous):

solution with the integral is ok,i will do that

OpenStudy (anonymous):

it is symmetric, so we do two times the region where y is positive to get rid of absolute value

OpenStudy (anonymous):

just was confuse in the formation

OpenStudy (anonymous):

thats right too

OpenStudy (anonymous):

looks like a heart shaped region

OpenStudy (anonymous):

i have a maple graph if you want to see it

OpenStudy (anonymous):

correct, i had drawn that too :)

OpenStudy (anonymous):

the first one is parabola, the second one had two loops passing through origin

OpenStudy (anonymous):

yep that's how mine looks too.

OpenStudy (anonymous):

let me look at the integral you formed

OpenStudy (anonymous):

ok